Science Report – November 4th

Crew 214 Science Report 4 November 2019
SOL: 8
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: The fifth scientific EVA for the mission saw us returning to Tank Wash #2 and collecting 28 additional samples for the micrometeorite search. This is 40% of the planned samples.

Green Hab: Following the commencement of planting within the green hab, access through the tunnel between the main hab and the green hab has been opened. The plants were watered with approximately 3.5 litres of water, with a remaining 4.5 litres remaining in the watering can. The plant soil appeared fairly moist, so it is suggested crew refrain from over watering the new plants.

Science Dome: Samples returned from the morning’s EVA have now been processed within the science dome and limited observations of material recovered through the binocular microscope have been carried out.

Closed Loop Food Waste Study: The day’s waste for Sol 8 is: 51 grams; consisting of 159 calories, 3g fat, >1g saturated fat, 199mg sodium, 42g carbohydrates, 3g fibre, 13g sugar, 6g protein. Extra food supplies were received and added to the baseline stock database.

EVA: Preparations for an EVA to White Moon at the junction between Brahe Highway and Cow Dung Road have been carried out.

EVA Suit Maintenance: Images of helmet 5 have been forwarded to CapCom detailing the partial failure, the battery to Suit 10 has been examined and it is the suit itself that is faulty.

Additional Activities: Nil.

Copyright © The Mars Society. All rights reserved. | Main Site